OpenStudy (anonymous):

The following function is one-to-one. Find its inverse. Find the domain and range of f and f^-1 Please show all of your work. f(x)=8+x/x

OpenStudy (anonymous):

y=8/(x-1)

OpenStudy (anonymous):

Thanx, but can I get some of the work.

OpenStudy (anonymous):

yx=8+x x(y-1)=8 x=8/(y-1) interchange x and y

OpenStudy (anonymous):

Thanx!!